OpenStudy (anonymous):

Let f(x) = x2 - 3x - 7. Find f(-3). Question 1 options: 1) -7 2) 8 3) 11 4) 22

OpenStudy (anonymous):

1

OpenStudy (vijay):

11

OpenStudy (anonymous):

\[(-3)^2-3(-3)-7\] \[=9+9-7\] \[=11\]

OpenStudy (anonymous):

^ sorry i meant to write 11 not 1